OpenStudy (anonymous):

as I can solve this exercise? 2^x/ (4+4^x) dx

OpenStudy (loser66):

integral?

myininaya (myininaya):

if it an integral i would go with the substitution u=2^x first

OpenStudy (loser66):

why not u = 2+2^x?

myininaya (myininaya):

or \[\frac{1}{4 \ln(2)} \int\limits_{}^{}\frac{2^x \ln(2)}{1+(\frac{2^x}{2})^2} dx\] this will actually combine two sub into one by doing \[\tan(\theta)=\frac{2^x}{2}\]

OpenStudy (loser66):

If u = 2+2^x, then du = ln2 2^x dx, du/ln2 = 2^x dx so that \[\int \dfrac{1}{2ln2} \dfrac{1}{u}du\] \(= \dfrac{1}{2ln2}ln(2+2^x)+C\) Check my stuff, please

myininaya (myininaya):

what happened to the 4+4^x

OpenStudy (loser66):

oh yea, you are correct, professor. hahaha... learn a lot!! My bad.

myininaya (myininaya):

\[4+4^x=2 \cdot 2 + (2 \cdot 2)^x\] oh you tried to factor 2 out

myininaya (myininaya):

except that one 2 has a x power

OpenStudy (loser66):

yea, I see my mistake.